# Fino1: On the Transferability of Reasoning-Enhanced LLMs to Finance

### v1

@@ 3.2 Error Analysis
Figure 1: Error case 1.
Figure 2: Error case 2.
3.3 Future Direction
Based on the above findings, we observe that developing a powerful reasoning model for financial tasks presents unique challenges distinct from general-domain reasoning. Our analysis suggests three key areas for improvement in financial reasoning models:
(1) Enhancing models’ understanding of financial knowledge and terminology. While reasoning-enhanced models demonstrate strong performance in general domains, they often struggle with financial tasks due to a lack of domain-specific knowledge. Future work should explore strategies that incorporate financial corpora, structured financial reports, and regulatory documents to improve models’ comprehension of domain-specific terminology and contextual nuances.
(2) Improving models’ ability to handle multi-table reasoning and long financial contexts. Our results highlight that even large models exhibit significant performance drops when processing long financial reports and cross-referencing multiple tables. To address this, future reasoning stratigies should be focused on inferring logics among tables or menorization mechanisms to enhance long-context understanding.
(3) Deeper analysis of reasoning-enhanced strategies in models like GPT-o1 and DeepSeek-R1. Although GPT-o1 and DeepSeek-R1 are designed with advanced reasoning enhancements, they fail to show consistent improvements across financial tasks. GPT-o1 utilizes "chain-of-thought" prompting, generating intermediate reasoning steps to enhance structured problem-solving. However, this reasoning strategy does not lead to improvements in any of the financial benchmarks, as GPT-o1 underperforms compared to GPT-4o across all tasks, scoring 49.07 in FinQA and 56.0 in DM-Simplong. In contrast, DeepSeek-R1 introduces self-reflection capabilities through reinforcement learning, enabling the model to iteratively evaluate and refine its reasoning processes without supervised data. While this approach helps DeepSeek-R1 achieve strong performance in XBRL-Math (86.67), indicating improved structured numerical reasoning, it also leads to a drop in FinQA (65.13), suggesting challenges in financial text-based reasoning. These results highlight a fundamental difference between the two models — GPT-o1’s structured reasoning approach does not translate to domain-specific improvements, whereas DeepSeek-R1’s reinforcement learning strategy enhances structured data processing but remains limited in free form textual financial reasoning. Understanding these trade-offs is essential for designing reasoning strategies tailored to financial tasks.
Developing robust financial reasoning models requires both domain adaptation and architectural advancements to bridge the gap between general reasoning capabilities and financial-specific challenges. Future work should focus on integrating financially specialized pretraining, multi-table comprehension techniques, and long-context optimization to build models capable of handling complex financial documents and reasoning tasks effectively.
4 Related Work
4.1 General-Purpose and Reasoning-Enhanced LLMs
4.2 LLMs in Financial Applications
4.3 Reasoning Capabilities in Financial Tasks
Financial reasoning presents unique challenges that combine numerical computation, domain knowledge, and logical inference. Recent studies have explored LLMs’ capabilities in financial reasoning tasks, including numerical reasoning over financial documents Chen et al. (2021), mathematical problem-solving in financial contexts Srivastava et al. (2024), and structured data interpretation Liu et al. (2025). These works highlight the importance of both domain expertise and reasoning abilities in financial applications. Our work extends these research directions by systematically evaluating how reasoning enhancements in state-of-the-art LLMs translate to financial domain tasks, providing insights into the effectiveness of different reasoning strategies and identifying areas for improvement in financial reasoning capabilities.
5 Conclusion
In this study, we evaluated the performance of 16 recent reasoning-augmented LLMs and general-purpose LLMs on financial tasks, analyzing their capabilities in text understanding, numerical reasoning, and structured data processing. Our results indicate that while reasoning-enhanced models demonstrate strong performance in general tasks, they often fail to generalize effectively to financial reasoning due to insufficient domain adaptation. Additionally, models optimized for structured mathematical reasoning perform well in numerical tasks like XBRL-Math but struggle with financial text comprehension and long-context processing, highlighting the need for financial domain-specific adaptations. To address these limitations, we presented Fino1, a reasoning-enhanced LLMs based on Llama-3.1-8B-Instruct, leveraging reasoning paths curated from GPT-4o using only FinQA, a dataset with 5,000 financial reasoning samples. Our results demonstrate that, despite being trained on a single dataset, Fino1 achieved an overall performance improvement of approximately 10% across three financial benchmarks. This significant improvement underscores the potential of financial domain-specific data in enhancing LLMs’ financial reasoning capabilities. Future work can focus on enhancing financial knowledge adaptation, improving multi-table reasoning, and handling long-context financial documents. Additionally, refining domain-specific reasoning strategies through structured data reasoning, retrieval-augmented approaches, and domain-aware pretraining will be critical for bridging the gap between general reasoning capabilities and financial-specific challenges.

### v2

@@ Appendix C Error Analysis
C.1 Future Direction
Developing effective reasoning models for financial tasks poses distinct challenges compared to general-domain reasoning. Our findings suggest three key directions for improvement: (1) Enhancing financial knowledge and terminology understanding through domain-specific pretraining using financial corpora, structured reports, and regulatory documents; (2) Improving multi-table reasoning and long-context comprehension by developing strategies that better infer logic across tables and optimize memory mechanisms for processing lengthy financial documents; and (3) Refining reasoning-enhanced strategies, as current approaches like GPT-o1’s chain-of-thought prompting and DeepSeek-R1’s self-reflection via reinforcement learning yield mixed results—highlighting the need for more tailored reasoning strategies to address both structured numerical and unstructured textual financial reasoning.
@@ C.1 Future Direction
Appendix D Prompt
Prompt Template for DocMath-CompLong task You are a financial expert, you are supposed to answer the given question based on the provided financial document context. You need to first think through the problem step by step, documenting each necessary step. Then you are required to conclude your response with the final answer in your last sentence as Therefore, the answer is {final answer}’. The final answer should be a numeric value. ###Context <Full document here> ### Input <the question based on the document> Lets think step by step to answer the given question. ### Output
Prompt Template for DocMath-Simplong task You are a financial expert, you are supposed to answer the given question based on the provided financial document context. You need to first think through the problem step by step, documenting each necessary step. Then you are required to conclude your response with the final answer in your last sentence as Therefore, the answer is {final answer}’. The final answer should be a numeric value. ###Context <simple long context here> ### Input <the question based on the context> Lets think step by step to answer the given question. ### Output
Prompt Template for FinQA task Please answer the given financial question based on the context. Context: <context> Question: <the question based on the context> Answer:
Prompt Template for XBRL-Math task You are a financial expert tasked with carefully reading, analyzing, and answering the following eXtensible Business Reporting Language. Please follow the steps below: INPUT: Read the eXtensible Business Reporting Language (XBRL) question: <XBRL related question>, formula: <the formula related to this question>, and the explanation: <term explanation>. Provide only the final answer which is the numerical result of the calculation. For formulas like ROI, provide percentages. Never use the percent symbol in percentages. OUTPUT: